OpenStudy (anonymous):

For questions 16 and 17, a bowl contains 14 red chips, 9 blue chips, and 12 green chips. One chip is randomly drawn. 16. Find P(Blue or green) 17. Find the odds of drawing a red chip.

OpenStudy (anonymous):

P(blue or green ) = 21/35 or 3/5

OpenStudy (anonymous):

how did you get that? :)

OpenStudy (anonymous):

P(red) = 14/35 or 2/5

OpenStudy (anonymous):

total number of chips = 35 total blue + green = 9 + 12 = 21 probability = 21/35

OpenStudy (anonymous):

similarly for red chips (14 red)

OpenStudy (anonymous):

Oh okay thank you! Would you so happen to know what a stem-and-leaf plot is?

OpenStudy (anonymous):

i think its something like this suppose you have set of numbers 12 17 19 16 22 29 31 32 33 the stem is first digit in number so you set them out like this 12 6 7 9 22 9 31 2 3

OpenStudy (anonymous):

not sure if thats exactly right

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

that link gives you all the info

OpenStudy (anonymous):

i) P(Blue or green) = Prob of blue + Prob of Green. **when the word OR is involved, it means that u have to add the probs, when the word AND is involved u gotta multiply the probs, .'. 9/35+12/35 = 21/35 or 0.6

OpenStudy (anonymous):

The prob that the chip chosen is red = Number of red chips/total chips Prob (red)= 14/35 = 2/5 or 0.4